# On the best approximation by finite Gaussian mixtures

### v1

@@ A.1 Distances between probability measures
  • $\chi^{2}$-divergence $\chi^{2}(P\|Q)=\int(\frac{(p-q)^{2}}{q}\mathrm{d}\mu$;
@@ A.3 Orthogonal polynomials
Appendix B Alternative proofs of lower bounds via orthogonal expansion
@@ B.1 Lower bound for Gaussian distributions
B.2 Lower bound for uniform distributions
We give an alternative proof of the lower bound (24) for approximating uniform mixing distribution. The framework resembles the approach in Section 4.1, while we consider the $\chi^{2}$ divergence instead of working with the $\mathsf{TV}$ distance directly. First, we show that it suffices to derive the same lower bound (within a constant in the exponent) under the $\chi^{2}$-divergence. Next, we relate the $\chi^{2}$ variational lower bound to the spectrum of some weighted moment matrix, and then apply orthogonal expansion to bound the corresponding coefficient matrix as carried out in Proposition 3.
@@ B.2 Lower bound for uniform distributions
Since $P,Q_{a}\in{\mathcal{P}}^{\textsf{Bdd}}_{a}$, by Lemmas 6 and 9,
@@ C.1 Proofs of main theorems
Pick any $d\in\{\mathsf{TV},H,\mathsf{KL},\chi^{2}\}$ and denote $m=m^{\star}(\epsilon,{\mathcal{P}}^{\textsf{Bdd}}_{M},d)$. Combining the result in Theorems 3 and 6, we have that $\log\frac{1}{\epsilon}\asymp m\log\frac{m}{M^{2}}$ when $m\geq CM^{2}$ for some constant $C$, and $\log\frac{1}{M\epsilon}\lesssim\frac{m^{2}}{M^{2}}\lesssim\log\frac{1}{\epsilon}$ when $3\sqrt{\kappa}M\leq m\leq CM^{2}$. Also, the condition $M\leq\epsilon^{-c_{1}},c_{1}\in(0,1)$ implies that $\log\frac{1}{M\epsilon}\asymp\log\frac{1}{\epsilon}$. Following (54), we can rewrite the result as
@@
Consider any $d\in\{\mathsf{TV},H,\mathsf{KL},\chi^{2}\}$. Denote $m=m^{\star}(\epsilon,{\mathcal{P}}_{\alpha}(\beta),d)$.
First we prove the lower bound. By Theorem 5,
$\frac{1}{\epsilon^{2}\beta^{\frac{4}{2+\alpha}}m^{\frac{2\alpha}{2+\alpha}}}\log\frac{1}{\epsilon^{2}\beta^{\frac{4}{2+\alpha}}m^{\frac{2\alpha}{2+\alpha}}}=O_{\alpha}\left(\frac{1}{\beta^{2}\epsilon^{2}}\right),$
which implies that $m=\Omega_{\alpha}\left[\beta\left(\log\frac{1}{\beta^{2}\epsilon^{2}}\right)^{\frac{2+\alpha}{2\alpha}}\right]$. By the relation $\beta\leq\epsilon^{-c_{1}}$ for $c_{1}<1$, we have $m=\Omega_{\alpha}\left[\beta\left(\log\frac{1}{\epsilon^{2}}\right)^{\frac{2+\alpha}{2\alpha}}\right]$.
@@
$\mathsf{TV}\left(f_{Q},f_{P}\right)$ $\overset{\rm(a)}{\leq}P(I_{t})\mathsf{TV}\left(f_{Q},f_{P_{t}}\right)+P(I_{t}^{c})\mathsf{TV}\left(f_{Q},f_{P_{t}^{c}}\right)$
$\overset{\rm(b)}{\leq}\sqrt{\frac{1}{2}\chi^{2}\left(f_{Q}\|f_{P_{t}}\right)}+P(I_{t}^{c}),$
@@
${\mathcal{E}}^{\star}(m,P,\mathsf{TV})\overset{\rm(a)}{\geq}\sup_{\delta>0}\frac{\pi\min_{0\leq\theta\leq 2\pi}g^{\mathsf{wrap}}(\theta)}{(m+1)\exp(m^{2}\delta^{2}/2)}\overset{\rm(b)}{\geq}\frac{\pi h_{\alpha}\left(\frac{3\pi}{\delta}\right)}{2m\delta\exp(m^{2}\delta^{2}/2)},$
@@
where the last inequality follows from the Poisson tail bound $\mathbb{P}[X\geq L]\leq e^{-4d^{2}R^{2}}\left(\frac{4ed^{2}R^{2}}{L}\right)^{L}$ for $X\sim\operatorname{Poisson}(4d^{2}R^{2})$ and $L>4d^{2}R^{2}$ [Ver18, Exercise 2.3.3]. ∎
@@
$K^{2}\tilde{L}$ $\overset{\rm(a)}{\geq}K^{2}{\left\lfloor{\frac{L}{3K}}\right\rfloor}\geq{\left\lfloor{\frac{6\kappa_{d}M^{2}}{L}}\right\rfloor}^{2}{\left\lfloor{\frac{L^{2}}{18\kappa_{d}M^{2}}}\right\rfloor}\overset{\rm(b)}{\geq}\left(\frac{6}{7}\frac{6\kappa_{d}M^{2}}{L}\right)^{2}\left(\frac{6}{7}\frac{L^{2}}{18\kappa_{d}M^{2}}\right)\geq\kappa_{d}M^{2},$
@@
Then, we extend the wrapped density approach to multi-dimensional case. Given a density function $g$ supported on a subset of $\mathbb{R}^{d}$, define the corresponding wrapped density $g^{\mathsf{wrap}}$ on $[-\pi,\pi]^{d}$ as
@@
Let $L={\left\lceil{(m+1)^{1/d}}\right\rceil}$ satisfying $m+1\leq L^{d}\leq 2^{d}(L-1)^{d}\leq 2^{d}m$. Let $X\sim P=\mathrm{Unif}([-M/\sqrt{d},M/\sqrt{d}]^{d})$ and $\delta=\frac{\pi\sqrt{d}}{M}$. Then, ${\mathbf{A}}_{L}(\delta X)=I_{L^{d}}$ and thus the conclusion follows from (60). ∎
@@
$\overset{\rm(a)}{\leq}\frac{2}{P(A)}\left[\exp\left(-\frac{\log\kappa_{d}}{42\kappa_{d}}\frac{\tilde{m}^{2}}{t^{2}}\right)+\exp\left(-\frac{1}{2}\left(\frac{t}{\sigma}-\sqrt{d}\right)^{2}\right)\right]$

### v2

@@ A.1 Distances between probability measures
  • $\chi^{2}$-divergence $\chi^{2}(P\|Q)=\int\frac{(p-q)^{2}}{q}\mathrm{d}\mu$;
@@ A.2 Gauss quadrature
Algorithm 2 implements the Gauss quadrature rule using the Golub-Welsch method [GW69], with the moment vector as input. For further implementation details on truncated Gaussian distributions, see also [Bur14, Sec. 3.10].
Algorithm 2 Quadrature Rule
  Input: A valid moment vector ${\mathbf{m}}=(m_{1},\ldots,m_{2k-1})$
  Output: A $k$-atomic distribution $P=\sum_{i=1}^{k}w_{k}\delta_{x_{k}}$.
  Define the following degree-$k$ polynomial $P(x)$:
$P(x)=\det\begin{bmatrix}1&m_{1}&\cdots&m_{k}\\ \vdots&\vdots&\ddots&\vdots\\ m_{k-1}&m_{k}&\cdots&m_{2k-1}\\ 1&x&\cdots&x^{k}\end{bmatrix}$
  Let the nodes $\{x_{1},\ldots,x_{k}\}$ be the roots of the polynomial $P(x)$.
  Let the weights $w=\{w_{1},\ldots,w_{k}\}$ be defined as:
$w=\begin{bmatrix}1&1&\cdots&1\\ x_{1}&x_{2}&\cdots&x_{k}\\ \vdots&\vdots&\ddots&\vdots\\ x_{1}^{k-1}&x_{2}^{k-1}&\cdots&x_{k}^{k-1}\end{bmatrix}^{-1}\begin{bmatrix}1\\ m_{1}\\ \vdots\\ m_{k-1}\end{bmatrix}$
@@ A.3 Orthogonal polynomials
Appendix B Proofs of lower bounds via orthogonal expansion
@@ B.1 Lower bound for Gaussian distributions
B.2 Completing the proof of Theorem 7
We prove the lower bound (26) for approximating uniform mixing distribution, and thereby complete the proof of Theorem 7. The framework resembles the approach in Section 4.1, while we consider the $\chi^{2}$ divergence instead of working with the $\mathsf{TV}$ distance directly. First, we show that it suffices to derive the same lower bound (within a constant in the exponent) under the $\chi^{2}$-divergence. Next, we relate the $\chi^{2}$ variational lower bound to the spectrum of some weighted moment matrix, and then apply orthogonal expansion to bound the corresponding coefficient matrix as carried out in Proposition 3.
@@ B.2 Completing the proof of Theorem 7
Since $P,Q_{a}\in{\mathcal{P}}^{\textsf{Bdd}}_{a}$, by Lemmas 4 and 7,
@@
B.3 Proof of Proposition 4
Lemma 8.
For any distribution $P$ on $\mathbb{R}$, under the notations above,
$\|{\mathbf{R}}_{m}\|_{F}^{2}\leq\sum_{n=0}^{m}2^{2n}\kappa_{n}^{2}.$
Proof.
It follows that
[…]
$\int p_{j}(x)p_{k}(x)w(x)\mathrm{d}x=h_{j}{\mathbf{1}{\left\{{j=k}\right\}}}.$
Lemma 9.
Now, with Lemmas 8 and 9, we are ready to show the lower bound for ${\mathcal{P}}^{\textsf{Bdd}}_{M}$:
Proof of Proposition 4.
Let $b\in(0,\pi)$ and $\gamma=\sin\frac{b}{2}$. Recall that $\tilde{P}$ is defined as the distribution supported on $[-b,b]$ with a density function
Let $\{\Psi_{n}\}$ and $\{\tilde{\varPhi}_{n}\}$ be the sets of monic orthogonal polynomials on the unit circle associated with $Q$ and $\tilde{P}$ as defined in (12), and $\psi_{n}=\kappa_{n}\Psi_{n}$, $\tilde{\varphi}_{n}=\tilde{\kappa}_{n}\tilde{\varPhi}_{n}$ be the orthonormal polynomials, respectively. By definition, $\tilde{\varPhi}_{n}(z)=(-1)^{n}\Psi_{n}(-z)$ and $\tilde{\kappa}_{n}=\kappa_{n}$. Let $\mathbf{\tilde{R}}_{m}$ be the coefficient matrix of $\{\tilde{\varPhi}_{n}\}$ of order $m+1$. By Lemmas 8 and 9, we obtain that
$r_{n}(x)=x-\frac{1}{4r_{n-1}(x)},\quad n\in\mathbb{N}.$
$\|\tilde{{\mathbf{R}}}_{m}\|_{F}^{2}\leq\sum_{n=0}^{m}2^{2n+1}\gamma^{-2n}=2\frac{(2/\gamma)^{2m+2}-1}{(2/\gamma)^{2}-1}.$ (59)
We conclude the proof of (26) by plugging in the choice of $b$. ∎
@@ C.1 Proofs of main theorems
For $d$ satisfying Assumption 1, denote $m=m^{\star}(\epsilon,{\mathcal{P}}^{\textsf{Bdd}}_{M},d)$. Combining the result in Theorems 3 and 7, we have that $\log\frac{1}{\epsilon}\asymp m\log\frac{m}{M^{2}}$ when $m\geq CM^{2}$ for some constant $C$, and $\log\frac{1}{M^{2}\epsilon}\lesssim\frac{m^{2}}{M^{2}}\lesssim\log\frac{1}{\epsilon}$ when $3\sqrt{\kappa}M\leq m\leq CM^{2}$. Also, the condition $M\leq\epsilon^{-c_{1}},c_{1}\in(0,\frac{1}{2})$ implies that $\log\frac{1}{M^{2}\epsilon}\asymp\log\frac{1}{\epsilon}$. Following (60), we can rewrite the result as
@@
For $d$ satisfying Assumption 1, denote $m=m^{\star}(\epsilon,{\mathcal{P}}_{\alpha}(\beta),d)$. First we prove the lower bound. By Theorem 5 and Lemma 4,
$\frac{1}{\epsilon\beta^{\frac{4}{2+\alpha}}m^{\frac{2\alpha}{2+\alpha}}}\log\frac{1}{\epsilon\beta^{\frac{4}{2+\alpha}}m^{\frac{2\alpha}{2+\alpha}}}=O_{\alpha}\left(\frac{1}{\beta^{2}\epsilon}\right),$
which implies that $m=\Omega_{\alpha}\left[\beta\left(\log\frac{1}{\beta^{2}\epsilon}\right)^{\frac{2+\alpha}{2\alpha}}\right]$. By the relation $\beta\leq\epsilon^{-c_{1}}$ for $c_{1}\in(0,\frac{1}{2})$, we have $m=\Omega_{\alpha}\left[\beta\left(\log\frac{1}{\epsilon}\right)^{\frac{2+\alpha}{2\alpha}}\right]$.
@@
$\mathsf{TV}\left(f_{Q},f_{P}\right)$ $\overset{\mathrm{(a)}}{\leq}P(I_{t})\mathsf{TV}\left(f_{Q},f_{P_{t}}\right)+P(I_{t}^{c})\mathsf{TV}\left(f_{Q},f_{P_{t}^{c}}\right)$
$\overset{\mathrm{(b)}}{\leq}\sqrt{\frac{1}{2}\chi^{2}\left(f_{Q}\|f_{P_{t}}\right)}+P(I_{t}^{c}),$
@@
${\mathcal{E}}^{\star}(m,P,\mathsf{TV})\overset{\mathrm{(a)}}{\geq}\sup_{\delta>0}\frac{\pi\min_{0\leq\theta\leq 2\pi}g^{\mathsf{wrap}}(\theta)}{(m+1)\exp(m^{2}\delta^{2}/2)}\overset{\mathrm{(b)}}{\geq}\frac{\pi h_{\alpha}\left(\frac{3\pi}{\delta}\right)}{2m\delta\exp(m^{2}\delta^{2}/2)},$
@@
where the last inequality follows from the Poisson tail bound $\mathbb{P}[X\geq L]\leq e^{-4d^{2}R^{2}}\left(\frac{4ed^{2}R^{2}}{L}\right)^{L}$ for $X\sim\operatorname{Poisson}(4d^{2}R^{2})$ and $L>4d^{2}R^{2}$ [MU05, Theorem 4.4]. ∎
@@
$K^{2}\tilde{L}$ $\overset{\mathrm{(a)}}{\geq}K^{2}{\left\lfloor{\frac{L}{3K}}\right\rfloor}\geq{\left\lfloor{\frac{6\kappa_{d}M^{2}}{L}}\right\rfloor}^{2}{\left\lfloor{\frac{L^{2}}{18\kappa_{d}M^{2}}}\right\rfloor}\overset{\mathrm{(b)}}{\geq}\left(\frac{6}{7}\frac{6\kappa_{d}M^{2}}{L}\right)^{2}\left(\frac{6}{7}\frac{L^{2}}{18\kappa_{d}M^{2}}\right)\geq\kappa_{d}M^{2},$
@@
Then, we extend the wrapped density approach and the orthogonal expansion approach described in Section 4.1.2 to multi-dimensional case.
Wrapped density.
Given a density function $g$ supported on a subset of $\mathbb{R}^{d}$, define the corresponding wrapped density $g^{\mathsf{wrap}}$ on $[-\pi,\pi]^{d}$ as
@@
Orthogonal expansion.
We recall some notations in Section 4.1.2. Let $\{\varphi_{n}\}$ be the orthonormal polynomials on the unit circle associated with $Y=\delta X$, where $X\sim P$ is a one-dimensional random variable and $\delta>0$. Denote ${\mathbf{R}}_{L-1,\delta}=(R_{jk})_{j,k=0}^{L-1}$ as the associated coefficient matrix of $\{\varphi_{k}\}$ of order $L$. For each multi-index $\boldsymbol{\ell}=(\ell_{0},\ldots,\ell_{d-1})\in\mathbb{Z}_{\geq 0}^{d}$, we define the $\boldsymbol{\ell}^{\mathrm{th}}$ multivariate orthogonal polynomial as
$\varphi_{\boldsymbol{\ell}}({\mathbf{w}})=\prod_{j=0}^{d-1}\varphi_{\ell_{j}}(w_{j}),\quad w\in\mathbb{C}^{d}.$
Then, $\{\varphi_{\boldsymbol{\ell}}({\mathbf{w}})\}$ is a basis of $d$-variate polynomial space. Let ${\mathbf{Y}}=(Y_{0},...,Y_{d-1})$ be $d$ i.i.d. copies of $Y$. By the orthogonality and independence,
$\mathbb{E}\left[\varphi_{\boldsymbol{\ell}}(e^{i{\mathbf{Y}}})\overline{\varphi_{\boldsymbol{\ell^{\prime}}}(e^{i{\mathbf{Y}}})}\right]=\prod_{j=0}^{d-1}\mathbb{E}\left[\varphi_{\ell_{j}}(e^{iY_{j}})\overline{\varphi_{\ell_{j}^{\prime}}(e^{iY_{j}})}\right]={\mathbf{1}{\left\{{\boldsymbol{\ell}=\boldsymbol{\ell}^{\prime}}\right\}}}.$ (67)
The next proposition lower bounds the minimum eigenvalue of multilevel Toeplitz matrices as an extension of Proposition 3.
Proposition 10.
Let $X\sim P$, $\delta>0$, and ${\mathbf{Y}}=(Y_{0},...,Y_{d-1})$, where each $Y_{j}$ is an i.i.d. copy of $Y=\delta X$. Then, with the above notations,
$\lambda_{\min}({\mathbf{A}}_{L}({\mathbf{Y}}))\geq\frac{1}{\|{\mathbf{R}}_{L-1,\delta}\|_{F}^{2d}}.$
Proof.
Let ${\mathbf{w}}=(w_{0},\ldots,w_{d-1})^{\top}\in\mathbb{C}^{d}$ with $w_{j}=e^{iy_{j}}$. Denote $\boldsymbol{\ell}_{k}=(\ell_{k,0},\ldots,\ell_{k,d-1})^{\top}\in\{0,1,\dots,L-1\}^{d}$ as the $L$-base representation vector for $k=\sum_{j=0}^{d-1}\ell_{k,j}L^{j}\in\{0,1,\dots,L^{d}-1\}$. For any ${\mathbf{x}}=(x_{0},\ldots,x_{L^{d}-1})^{\top}\in\mathbb{C}^{L^{d}}\backslash\{\mathbf{0}\}$, let $\pi({\mathbf{w}})=\sum_{k=0}^{L^{d}-1}x_{k}{\mathbf{w}}^{\boldsymbol{\ell}_{k}}$ with ${\mathbf{w}}^{\boldsymbol{\ell}_{k}}\triangleq\prod_{j=0}^{d-1}w_{j}^{\ell_{k,j}}$. Expand $\pi({\mathbf{w}})$ under the orthogonal basis $\{\varphi_{\boldsymbol{\ell}_{k}}\}$ as
$\pi({\mathbf{w}})=\sum_{k=0}^{L^{d}-1}x_{k}{\mathbf{w}}^{\boldsymbol{\ell}_{k}}=\sum_{k=0}^{L^{d}-1}c_{k}\varphi_{\boldsymbol{\ell}_{k}}({\mathbf{w}}).$
Letting ${\mathbf{c}}=(c_{0},\ldots,c_{L^{d}-1})^{\top}$ and ${\mathbf{A}}\equiv{\mathbf{A}}_{L}({\mathbf{Y}})$, the orthogonality of $\{\varphi_{\boldsymbol{\ell}_{k}}\}$ in (67) implies that
$\|{\mathbf{c}}\|^{2}=\mathbb{E}\left[|\pi(e^{i{\mathbf{Y}}})|^{2}\right]={\mathbf{x}}^{\star}{\mathbf{A}}{\mathbf{x}},$
where the last equality follows from the fact that $A_{kk^{\prime}}=\mathbb{E}\exp(i(\boldsymbol{\ell}_{k}-\boldsymbol{\ell}_{k^{\prime}})^{\top}{\mathbf{Y}})$.
Let ${\mathbf{R}}\equiv{\mathbf{R}}_{L-1,\delta}$ be the coefficient matrix of $\{\varphi_{k}\}$ of order $L$. Note that
$\pi({\mathbf{w}})$ $=\sum_{k=0}^{L^{d}-1}c_{k}\varphi_{\boldsymbol{\ell}_{k}}({\mathbf{w}})$
$=\sum_{k=0}^{L^{d}-1}c_{k}\left[\prod_{j=0}^{d-1}\left(\sum_{t=0}^{\ell_{k,j}}R_{\ell_{k,j}t}w_{j}^{t}\right)\right]$
$=\sum_{k^{\prime}=0}^{L^{d}-1}\left[\sum_{k:\boldsymbol{\ell}_{k^{\prime}}\leq\boldsymbol{\ell}_{k}}c_{k}\left(\prod_{j=0}^{d-1}R_{\ell_{k,j}\ell_{k^{\prime},j}}\right)\right]{\mathbf{w}}^{\boldsymbol{\ell}_{k^{\prime}}},$
where $\boldsymbol{\ell}_{k^{\prime}}\leq\boldsymbol{\ell}_{k}$ denotes $\ell_{k^{\prime},j}\leq\ell_{k,j}$ for all $j=0,\ldots,d-1$. By comparing the coefficients, we have that ${\mathbf{x}}^{\top}={\mathbf{c}}^{\top}{\mathbf{S}}$, where ${\mathbf{S}}=(S_{kk^{\prime}})_{k,k^{\prime}=0}^{L^{d}-1}\in\mathbb{R}^{L^{d}\times L^{d}}$ is defined as $S_{kk^{\prime}}=\prod_{j=0}^{d-1}R_{\ell_{k,j}\ell_{k^{\prime},j}}$ if $\boldsymbol{\ell}_{k^{\prime}}\leq\boldsymbol{\ell}_{k}$, and $S_{kk^{\prime}}=0$ otherwise. Then, by definition,
$\|{\mathbf{S}}\|_{F}^{2}$ $=\sum_{k=0}^{L^{d}-1}\sum_{k^{\prime}:\boldsymbol{\ell}_{k^{\prime}}\leq\boldsymbol{\ell}_{k}}\prod_{j=0}^{d-1}R_{\ell_{k,j}\ell_{k^{\prime},j}}^{2}$
$=\sum_{0\leq\ell^{\prime}_{0}\leq\ell_{0}\leq L-1}\cdots\sum_{0\leq\ell^{\prime}_{d-1}\leq\ell_{d-1}\leq L-1}\prod_{j=0}^{d-1}R_{\ell_{j}\ell^{\prime}_{j}}^{2}$
$=\prod_{j=0}^{d-1}\left(\sum_{0\leq\ell_{j}^{\prime}\leq\ell_{j}\leq L-1}R_{\ell_{j}\ell^{\prime}_{j}}^{2}\right)=\|{\mathbf{R}}\|_{F}^{2d}.$
Hence, $\|{\mathbf{x}}\|\leq\|{\mathbf{c}}\|\|{\mathbf{S}}\|\leq\|{\mathbf{c}}\|\|{\mathbf{S}}\|_{F}=\|{\mathbf{c}}\|\|{\mathbf{R}}\|_{F}^{d}$. Finally, applying (19), we have
$\lambda_{\min}({\mathbf{A}})=\min_{{\mathbf{x}}\neq\mathbf{0}}\frac{{\mathbf{x}}^{\star}{\mathbf{A}}{\mathbf{x}}}{\|{\mathbf{x}}\|^{2}}\geq\min_{{\mathbf{x}}\neq\mathbf{0}}\frac{\|{\mathbf{c}}\|^{2}}{\|{\mathbf{c}}\|^{2}\|{\mathbf{R}}\|_{F}^{2d}}=\frac{1}{\|{\mathbf{R}}\|_{F}^{2d}}.\qed$
@@
Let $L={\left\lceil{(m+1)^{1/d}}\right\rceil}$ satisfying $m+1\leq L^{d}\leq 2^{d}(L-1)^{d}\leq 2^{d}m$. Let $X\sim P=\mathrm{Unif}([-M/\sqrt{d},M/\sqrt{d}]^{d})$ and $\delta=\frac{\pi\sqrt{d}}{M}$. Then, ${\mathbf{A}}_{L}(\delta X)=I_{L^{d}}$ and thus the conclusion (37) follows from (66).
Next we prove (38). Assume that $L\geq\frac{eM^{2}}{2d}$. Set $b={\sqrt{\frac{M^{2}}{Ld}\log\frac{2Ld}{eM^{2}}}}\in(0,\frac{\sqrt{2}}{e})$ and $\delta=b\frac{\sqrt{d}}{M}$. Let $Y\sim\tilde{P}$ be as defined in (29) that is supported on $[-b,b]$, and denote $P$ as the distribution of $\frac{Y}{\delta}$ supported on $[-\frac{M}{\sqrt{d}},\frac{M}{\sqrt{d}}]$. Consider the test distribution $P^{\otimes d}\in{\mathcal{P}}^{\textsf{Bdd}}_{M,d}$. Let ${\mathbf{R}}$ be the order-$L$ associated coefficient matrix of the orthogonal system with $\tilde{P}$. Similar to (59), we have $\|{\mathbf{R}}\|_{F}^{2}\leq\exp\left(O(L\log\frac{1}{b})\right)$. Applying (66) and Proposition 10, we have
${\mathcal{E}}^{\star}(m,P,\mathsf{TV})$ $\geq\frac{1}{2L^{d}}\exp\left[-\Omega\left(d\delta^{2}L^{2}+dL\log\frac{1}{b}\right)\right]$
$\geq\frac{1}{2^{d+1}m}\exp\left[-\Omega\left(dL\log\frac{dL}{M^{2}}\right)\right].$
The desired result follows. ∎
@@
$\overset{\mathrm{(a)}}{\leq}\frac{2}{P(A)}\left[\exp\left(-\frac{\log\kappa_{d}}{42\kappa_{d}}\frac{\tilde{m}^{2}}{t^{2}}\right)+\exp\left(-\frac{1}{2}\left(\frac{t}{\sigma}-\sqrt{d}\right)^{2}\right)\right]$